# Characteristics of Dark Current from an S-band RF Gun with Exchangeable Photocathode System

### v1

@@ V CONCLUSIONS
Despite the lack of quantitative convergence, the results above demonstrate the usefulness of simulations in the ability to characterize the effect of each emitting source on the screen pattern. In the screen images of the dark current the emission from the backplate and from the cathode insertion has been identified. This is clear due to the comparison of different experimental conditions (retraction and reinsertion of cathodes), and is also supported by the simulations, where different emission sources can be isolated, and shows some level of agreement with observation. For the CLARA gun, these observations lead to the following conclusions.
@@
Field emission from the cathode (flat surface and and the curved rim) leads to features observed in the central part of the screen, when the solenoid strength is set approximately to its nominal beam focussing setting. The central bright spot may not be due to ‘field emitters’ in the centre of the cathode, since simulations indicate field emission from the edge of the cathode may be focused to the centre of the screen.
There are several issues and uncertainties with the simulations that would require further work to evaluate further. There are various ways in which the electric and magnetic field in the simulations may differ from reality. At the emission surfaces where the particle energy is low, this may significantly affect the dynamics of the field emission. In simulation, meshing artefacts may affect the gun electromagnetic field shape, even for perfectly described geometries. Other effects which might be thought small may also be important; for example mutual coil excitation included in the modelling of combined solenoid and bucking solenoid field were observed to make a significant different to the simulated screen patterns. We do not have detailed in-situ 3-D measurements of the magnetic and electric fields, which may include other such small effects, to compare against the modelled fields.

### v3

@@ V CONCLUSIONS
Despite the lack of quantitative convergence, the results above demonstrate the usefulness of simulations in the ability to characterise the effect of each emitting source on the screen pattern. In the screen images of the dark current the emission from the backplate rim and from the cathode insertion has been identified. This is clear due to the comparison of different experimental conditions (retraction and reinsertion of cathodes), and is also supported by the simulations, where different emission sources can be isolated and the simulated screen features (rings, bright and dark regions) match the measured features in size and shape. For the CLARA gun, these observations lead to the following conclusions.
@@
Field emission from the cathode (flat surface and the curved rim) leads to features observed in the central part of the screen, when the solenoid strength is set approximately to its nominal beam focussing setting. The central bright spot may not be due to field emitters in the centre of the cathode, since simulations indicate field emission from the edge of the cathode may be focused to the centre of the screen.
There are several issues and uncertainties with the simulations that would require further work to evaluate. Our hypothesis on the effect of the cathode position or length on the dark current brightness has not yet been simulated in full. Futhermore there are various ways in which the electric and magnetic fields in the simulations may differ from reality. At the emission surfaces where the particle energy is low this may significantly affect the dynamics of the field emission. In simulation, meshing artefacts may affect the gun electromagnetic field shape, even for perfectly described geometries. Other effects which might be thought small may also be important; for example mutual coil excitation included in the modelling of combined solenoid and bucking solenoid field were observed to make a significant difference to the simulated screen patterns. We do not have detailed in-situ 3-D measurements of the magnetic and electric fields, which may include other such small effects, to compare against the modelled fields.